OpenStudy (anonymous):

Giving medals! Given: 3x + y = 1. Solve for y. y = -3x + 1 y = -3x - 1 y = 3x - 1

OpenStudy (anonymous):

set the equation = to y

OpenStudy (anonymous):

then solve it

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

You would start out by subtracting y from both sides

OpenStudy (anonymous):

You are supposed to help not just give the answers

OpenStudy (anonymous):

subtract Y from what!

OpenStudy (anonymous):

acually you would subtract the 3x @CookieMonsta22

OpenStudy (anonymous):

Oh yeah that would make it easier! My bad :/

OpenStudy (anonymous):

all you have to do is subtract 3x and since that is not one of the choices add -3x because its basically the same thing

OpenStudy (anonymous):

once you do that it would be y by itself

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Then you would simplify

OpenStudy (anonymous):

im terrible with this stuff. i got another one anyone want to help? Given: x + 2y = -6. Solve for y.

OpenStudy (anonymous):

the answer choices dont look like you would simplify

OpenStudy (anonymous):

that one you'd still set it = to y and get 2y=-x-6

OpenStudy (anonymous):

you would subtract the x first and then divide by 2

OpenStudy (anonymous):

Not for this question but if you were to solve for y then you would continue by simplifying.

OpenStudy (anonymous):

that is not one of the answers its all a fraction like y= x-6 over 2

OpenStudy (anonymous):

the answer is -6 -x over 2

OpenStudy (anonymous):

ohh well cool. thanks

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

\[y=\frac{ -x-6 }{ 2 }\]

OpenStudy (anonymous):

this ones a bit harder : Given: x + y = 6. If x {-15, -6, -1}, then which of the following sets of ordered pairs are solutions? {(-15, -21), (-6, -12), (-1, -7)} {(-15, 21), (-6, 12), (-1, 7)} {(-15, 21), (-6, 12), (-1, -7)}

OpenStudy (anonymous):

how did you get the pic on here?

OpenStudy (anonymous):

it has an equation option

OpenStudy (anonymous):

click equation by the chat and type it accordenly

OpenStudy (anonymous):

that one is easy just plug in the x's and solve for the y's @scoob

OpenStudy (anonymous):

ohhh thats cool good to know.

OpenStudy (anonymous):

i got it!

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

If the ordinate equals the abscissa and x {-3, 0, 12}, which of the following sets of ordered pairs satisfies the given conditions? {(3, -3), (0, 0), (-12, 12)} {(-3, 3), (0, 0), (12, -12)} {(-3, -3), (0, 0), (12, 12)}

OpenStudy (anonymous):



OpenStudy (anonymous):

isnt the abscissa the same as x?

OpenStudy (anonymous):

yes i would think so

OpenStudy (anonymous):